# Integrated population model reveals human and environment driven changes in Baltic ringed seal (Pusa hispida botnica) demography and behavior

### v1

@@ Acknowledgements
where $\tau$ is the time since the onset of the hunting season, $H(\tau)$ is the total size of the hunting bag at time $\tau$, $\tilde{n}_{i}(\tau)$ is the number (or density on ice) of seals of class $i$ that are still alive at time $\tau$, and $\varepsilon$ is a constant governing the magnitude of hunting relative to the population size, often called a timescale parameter (kuehn2015). If we assume that the number of hunting licenses is small, we can take the limit $\varepsilon\to 0$ to decouple the second ODE from the first. In other words, we assume that the dynamics of hunting and natural mortality unfold over different timescales, and approximate the joint dynamics of the population and the hunting bag as though the population was unaffected by hunting.
@@
The total rate of hunting is given by a sum of Equation S5 over all demographic classes,
@@
which is separable, and has the explicit time-dependent solution for the size of the hunting bag,
@@
To estimate the composition of the hunting bag, we first divide Equation S5 by Equation S6,
@@
Dividing both sides by the total size of the hunting bag and evaluating at a time corresponding to the length of the hunting season, $\tau^{*}$, yields an expression for the expected composition of the hunting bag
@@
Since $H^{\prime}(\tau)$ integrates to $H(\tau^{*})$ over the interval $[0,\tau^{*}]$, $h(\tau)$ is a valid probability density function defined in this interval, and represents the conditional probability that a randomly sampled seal from the hunting bag was hunted within the time interval $[\tau,\tau+d\tau]$. The composition of the hunting bag is therefore an expectation of $g_{i}(\tau)$ over the distribution $h(\tau)$:
@@
The relative probability that a seal of sex $s$ and age $a$ becomes hunted in Sweden either during the spring or the fall can be approximated by normalizing Equations S16 and S17, which results in:
@@
Even though we did not explicitly model bycatch, an expression analogous to Equations S18 and S19 can be obtained to approximate the relative probabilities of becoming bycaught. Denoting by $B_{s,a}$ the actual number of bycaught seals with sex $s$ and age $a$, and ignoring the effects of demographic stochasticity, $B_{s,a}$ can be written as
@@
where $D_{s,a}=(1-\phi_{s,a})n_{s,a}$ is the number of seals of sex $s$ and age $a$ that died due to natural causes, including bycatch. A comparison of Equation S21 with the observation model for bycaught samples in Equation (19) indicates that
@@
Appendix Appendix S4 Density dependence
@@
We assumed that the failure rate of pregnancies (e.g. due to abortions) is highest immediately after mating, and declines linearly to zero by parturition. Moreover, we assumed that this rate, denoted by $\theta$, increases exponentially with population size:
$\theta(N,\tau)=\theta^{\text{max}}_{0}e^{\theta_{1}N}(1-\tau).$ (S1)
@@
where the second equality follows from the fact that $p_{0}$ is independent of $N$. Substituting Equation S4 into Equation S3,
$p_{t-1}(N_{t-1},\tau)=b_{t}(N_{t-1})e^{\theta_{0}e^{\theta_{1}N_{t-1}}(1-(2-\tau)\tau)}$ (S5)
@@
$p_{t-1}(N_{t-1},\tau)=b_{t}(0)e^{\theta_{0}(1-e^{\theta_{1}N_{t-1}}(2-\tau)\tau)}$ (S6)
where subscripts are included to make explicit the fact that both the population size and reproductive rates depend on the year, $t$. Setting $\tau=1$ in Equation S6 yields an expression for the density dependent birth rate:
@@
It is useful to parametrize Equation S7 in terms of the carrying capacity $K$ instead of the parameter $\theta_{1}$ since information on historic population sizes of Baltic ringed seals is much more prevalent in the literature than information on the mechanisms of density dependent regulation. Evaluating Equation S7 at the carrying capacity, and rearranging the terms to isolate $\theta_{1}$, gives
@@
where we have assumed that $K$ is the carrying capacity for a ringed seal population in the absence of the effects of organochlorine contamination on reproduction. The value of $b(K)$ can be found from Equation (25) by setting $\lambda=1$:
@@
Appendix Appendix S5 Observation models for sampling bias in Swedish hunting during the spring
@@
Since no rules have been specified in 2015, we assumed that seals were sampled randomly from the hunting bag (i.e. $v_{s,a}=1,\ \forall s,a$).
@@
We assumed that any sampling bias would be identical between 2016 and 2019. Using hunting records from 2019, we modelled whether or not seal $i$ is sampled from the hunting bag using the Bernoulli distribution
@@
The probability $v_{s,a}$ that a seal with sex $s$ and age $a$ is sampled from the Swedish hunting bag during the spring was then calculated as
@@
Appendix Appendix S6 Observation models for incomplete data
Hunted or bycaught seal samples occasionally had missing age or sex information. We assumed that the age and sex information was missing at random from those samples. We therefore modelled samples with missing sex or age information using a multinomial distribution similar to those presented in Section 2.4.2, but with probability vectors given respectively by the age and sex composition of the true hunting bag and bycatch.
@@
Appendix Appendix S7 Implementation in Stan

### v2

@@ Acknowledgements
where $\tau$ is the time since the onset of the hunting season, $H(\tau)$ is the total harvest at time $\tau$, $\tilde{n}_{i}(\tau)$ is the number (or density on ice) of seals of class $i$ that are still alive at time $\tau$, and $\varepsilon$ is a constant governing the magnitude of hunting relative to the population size, often called a timescale parameter (kuehn2015). If we assume that the number of hunting licenses is small, we can take the limit $\varepsilon\to 0$ to decouple the second ODE from the first. In other words, we assume that the dynamics of hunting and natural mortality unfold over different timescales, and approximate the joint dynamics of the population and the harvests as though the population was unaffected by hunting.
@@
The total rate of hunting is given by a sum of Equation (S5) over all demographic classes,
@@
which is separable, and has the explicit time-dependent solution for the harvest totals,
@@
To estimate the demographic composition of the harvests, we first divide Equation (S5) by Equation (S6),
@@
Dividing both sides by the total harvests and evaluating at a time corresponding to the length of the hunting season, $\tau^{*}$, yields an expression for the expected demographic composition of the harvests
@@
Since $H^{\prime}(\tau)$ integrates to $H(\tau^{*})$ over the interval $[0,\tau^{*}]$, $h(\tau)$ is a valid probability density function defined in this interval, and represents the conditional probability that a randomly sampled seal from the harvests was hunted within the time interval $[\tau,\tau+d\tau]$. The composition of the harvest is therefore an expectation of $g_{i}(\tau)$ over the distribution $h(\tau)$:
@@
The relative probability that a seal of sex $s$ and age $a$ becomes hunted in Sweden either during the spring or the fall can be approximated by normalizing Equations (S16) and (S17), which results in:
@@
Even though we did not explicitly model bycatch, an expression analogous to Equations (S18) and (S19) can be obtained to approximate the relative probabilities of becoming bycaught. Denoting by $B_{s,a}$ the actual number of bycaught seals with sex $s$ and age $a$, and ignoring the effects of demographic stochasticity, $B_{s,a}$ can be written as
@@
where $D_{s,a}=(1-\phi_{s,a})n_{s,a}$ is the number of seals of sex $s$ and age $a$ that died due to natural causes, including bycatch. A comparison of Equation (S21) with the observation model for bycaught samples in Equation (19) indicates that
@@
Appendix Supplement 4 Density dependence
@@
Given reports of significant late-term pregnancy losses in other pinniped populations and the increased energetic demands of late-stage gestation (pitcher1998; mckenzie2005; stenson2016), we assumed the rate of pregnancy losses increases linearly from zero at conception to a maximum at parturition. Moreover, we assumed that this rate, denoted by $\theta$, increases exponentially with population size:
$\theta(N,\tau)=\theta^{\text{max}}_{0}e^{\theta_{1}N}\tau.$ (S1)
@@
where the second equality follows from the fact that $p_{0}$ is independent of $N$. Substituting Equation (S4) into Equation (S3),
$p_{t-1}(N_{t-1},\tau)=b_{t}(N_{t-1})e^{\theta_{0}e^{\theta_{1}N_{t-1}}(1-\tau^{2})}$ (S5)
@@
$p_{t-1}(N_{t-1},\tau)=b_{t}(0)e^{\theta_{0}(1-e^{\theta_{1}N_{t-1}}\tau^{2})}$ (S6)
where subscripts are included to make explicit the fact that both the population size and reproductive rates depend on the year, $t$. Setting $\tau=1$ in Equation (S6) yields an expression for the density dependent birth rate:
@@
It is useful to parametrize Equation (S7) in terms of the carrying capacity $K$ instead of the parameter $\theta_{1}$ since information on historic population sizes of Baltic ringed seals is much more prevalent in the literature than information on the mechanisms of density dependent regulation. Evaluating Equation (S7) at the carrying capacity, and rearranging the terms to isolate $\theta_{1}$, gives
@@
where we have assumed that $K$ is the carrying capacity for a ringed seal population in the absence of the effects of organochlorine contamination on reproduction. The value of $b(K)$ can be found by setting the net reproduction rate equal to one (caswell2001; ersalman2024):
@@
Since there is little evidence of late-term pregnancy losses in Baltic ringed seals, and boyd1991 has postulated that the influence of nutritional factors on pinniped reproduction is greatest at the point of implantation, we also considered a model where the rate of pregnancy losses was assumed to decrease linearly:
$\theta(N,\tau)=\theta^{\text{max}}_{0}e^{\theta_{1}N}(1-\tau).$ (S10)
With this assumption, the density dependent pregnancy rate is given by
$p_{t-1}(N_{t-1},\tau)=b_{t}(0)e^{\theta_{0}(1-e^{\theta_{1}N_{t-1}}(2-\tau)\tau)}.$ (S11)
Appendix Supplement 5 Observation models for sampling bias in Swedish hunting during the spring
@@
Since no rules have been specified in 2015, we assumed that seals were sampled randomly from the harvests (i.e. $v_{s,a}=1,\ \forall s,a$).
@@
We assumed that any sampling bias would be identical between 2016 and 2019. Using hunting records from 2019, we modelled whether or not seal $i$ is sampled from the harvests using the Bernoulli distribution
@@
The probability $v_{s,a}$ that a seal with sex $s$ and age $a$ is sampled from the Swedish harvest during the spring was then calculated as
@@
Appendix Supplement 6 Observation models for incomplete data
Hunted or bycaught seal samples occasionally had missing age or sex information. We assumed that the age and sex information was missing at random from those samples. We therefore modelled samples with missing sex or age information using a multinomial distribution similar to those presented in Section 2.4.2, but with probability vectors given respectively by the age and sex composition of the true harvests and bycatch.
@@
Appendix Supplement 7 Implementation in Stan